Please help!
Solve by elimination:
-4x+3y=-1
8x+6y=14
Thank you so much!! Trying to help my daughter with her homework and not remembering at all how to do these.

Answer by Alan3354(69443)

Solve by elimination:
-4x+3y=-1
8x+6y=14 Divide by 2
---------------
-4x+3y=-1
4x+3y = 7
----------- Add
0x + 6y = 6
y = 1
--------
4x+3y = 7
4x + 3 = 7
4x = 4
x = 1